[figure: 9.7 cm]

MAGNETOOPTICAL MICROMAGNETOMETRY OF ULTRATHIN CO WEDGES.

Z.Kurant^{a}, M.Tekielak^{a},M.Kisielewski^{a}, A.Maziewski^{a} A.Wawro^{b}, L.T.Baczewski^{b}

aInstitute of Physics, University of Białystok, ul. Lipowa 41, 15-424 Białystok, Poland

bInstitute of Physics Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warsaw, Poland.

Ultrathin cobalt wedges Au/Co(d < 3nm)/Au(111)/Mo/Al_{2}O_{3}, grown by MBE, were investigated by means of the polar Kerr effect. Remnant states, affected by a magnetic history of the sample, were investigated by optical polarizing microscope equipped with a high sensitivity CCD camera. Micromagnetometric study, i.e. magnetometry in micrometer scale, were performed. Drastic changes in domain structure were observed with decrease of Co layer thickness below the magnetization reorientation transition from easy axis into easy plane : (i) decrease of domain nucleation density during remagnetization process; (ii) alteration of wall fractal structure in the thickness region, where domain wall propagation mechanism dominates in the process of remagnetization. Also wall velocity dependence on applied field amplitude was measured. Similar changes in domain properties were observed during sample transition into superparamagnetic state. The investigations were compared with classical magnetooptical milimagnetometric study, i.e magnetometry in milimiter scale.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : A. Maziewski

Address for correspondence : A. Maziewski Lipowa 41 Street 15-424 Białystok

Email address : magnet@cksr.ac.bialystok.pl